40m QIL Cryo_Lab CTN SUS_Lab TCS_Lab OMC_Lab CRIME_Lab FEA ENG_Labs OptContFac Mariner WBEEShop
  40m Log  Not logged in ELOG logo
Entry  Thu Jun 4 20:54:40 2020, gautam, Update, SUS, MC1 Slow Bias issues 
    Reply  Thu Jun 4 21:32:00 2020, Koji, Update, SUS, MC1 Slow Bias issues 
       Reply  Thu Jun 25 15:11:00 2020, gautam, Update, SUS, MC1 coil driver resistance quartered MCF.pdf
          Reply  Sun Jun 28 15:30:52 2020, gautam, Update, SUS, MC1 sat-box de-lidded 20200628T144138.jpg
             Reply  Sun Jun 28 16:29:58 2020, rana, Update, SUS, MC1 sat-box de-lidded 
                Reply  Sun Jun 28 17:36:35 2020, gautam, Update, SUS, MC1 sat-box de-lidded 
                   Reply  Mon Jun 29 11:55:46 2020, gautam, Update, SUS, MC1 sat-box de-lidded SatBoxSwap.jpg
                      Reply  Mon Jun 29 20:30:53 2020, Koji, Update, SUS, MC1 sat-box de-lidded 
Message ID: 15376     Entry time: Thu Jun 4 20:54:40 2020     Reply to this: 15377
Author: gautam 
Type: Update 
Category: SUS 
Subject: MC1 Slow Bias issues 

Summary:

I found that there is an issue with the MC1 slow bias voltages. 

Details:

I usually offload the DC part of the output voltage from the WFS servos to the slow bias voltage sliders, so as to preserve maximum actuation range from the fast system. However, today, I found that this servo wasn't working well at all. So I dug a little deeper. Looking at the EPICS database records:

  • The user-facing channels are "PIT" and "YAW" bias voltages.
  • These are converted to voltages to be sent to individual coils by some calc channels in the EPICS database record. So, for example, the voltage to be sent to the "UL" coil (Upper Left, as viewed from the AR side of the optic), is A+B, where A is the "PIT" voltage and B is the "YAW" voltage. Similar combinations of A and B are used for the other 3 face coils.
  • The problem is obvious - if either A or B > 5V, then the requested voltage to be sent to the UL coil is > 10 V, while the Acromag DACs can put out a maximum of 10 V
  • As it happens, with the IFO currently aligned, MC1 is the only optic which faces this problem. 
  • Why has this not been an issue before? In fact, looking at some old data, the "PIT" and "YAW" bias voltages to MC1 were both ~1-2 V in 2018. But I confirmed that something in the region of ~5 V is required from each of the "PIT" and "YAW" channels to bring the MCREFL spot back to the center of the camera, so something has changed the DC alignment of MC1, maybe an earthquake or something? Anyway, with these settings, 2/4 coils are basically saturated, and so we can only move the optic diagonally. 😢 
  • Other coils that have  requested output voltages > 5V (so more than half the range of the DAC) include MC2 LL (5.2V), and ETMX LL and LR (5.5 and 5.8 V respectively).
  • Either a factor of 0.5 should be included in all the EPICS database records, or else, we should make the "PIT" and "YAW" sliders range only from -5 to +5 V, so that this kind of misleading info isn't wasting time.
ELOG V3.1.3-